OpenStudy (anonymous):

Given vectors u= -7i + 4j and v= 8i + 9j, find 5u - 9v in terms of unit vectors i and j.

OpenStudy (xapproachesinfinity):

any thoughts

OpenStudy (anonymous):

i got -60i - 108j

OpenStudy (xapproachesinfinity):

what did you do

OpenStudy (xapproachesinfinity):

i can't say you are right or wrong unless i do it

OpenStudy (anonymous):

I got both the numbers but in my choices i didn't know which one to choose since I honestly just winged it. :( my choices are -104i - 108j -107i -61j -66i -108j - 60 - 108j

OpenStudy (xapproachesinfinity):

seems you got the last choice

OpenStudy (xapproachesinfinity):

D

OpenStudy (freckles):

do you know how to find 5u?

OpenStudy (freckles):

c<a,b>=<ca,cb> or the other notation (in terms of the unit vectors) c(ai+bj)=(ca)i+(cb)j

OpenStudy (anonymous):

I first did 5(-7 + 4j) - 9(8i +9j) but I realized none of my answers were in the choices so I switched 9 and 5.

OpenStudy (amity):

Correct answer is-107i-61j.. 5(-7i+4j)-9(8i+9j)=(-35-72)i +j(-81+20)=-107i-61j

OpenStudy (amity):

Correct answer is-107i-61j.. 5(-7i+4j)-9(8i+9j)=(-35-72)i +j(-81+20)=-107i-61j

OpenStudy (xapproachesinfinity):

what you did in the beginning is correct

OpenStudy (xapproachesinfinity):

5u=5(-7i+4j)=-35i+20j

OpenStudy (xapproachesinfinity):

-9v=-9(8i+9j)=-72i-81j

OpenStudy (anonymous):

yes i got that

OpenStudy (xapproachesinfinity):

combine them 5u-9v=-35i+20j-72i-81j=-107i-61j

OpenStudy (xapproachesinfinity):

apparently that's the second choice

OpenStudy (anonymous):

ooh i forgot i had to combine them. i was seriously just adding 35 and 20. thank you!

OpenStudy (xapproachesinfinity):

no problem